# VMem: Consistent Interactive Video Scene Generation with Surfel-Indexed View Memory

### v2

@@ 3.2 Novel View Generator
Figure 5: Generalization to long sequences with revisitations. We compare scene generation with and without our proposed memory module (VMem). For each sequence, an input image is shown on the left, followed by generated frames at selected frame indices. Our method (top row in each pair) maintains high spatial consistency across revisited viewpoints, while the baseline without memory (bottom row) suffers from visual drift and structural degradation over time.
We use a camera-conditioned image-set generator $\psi$ to generate new views. The generator $\psi$ takes $K$ retrieved reference views $\mathcal{V}^{*}=\{v_{t}=(\mathbf{x}_{t},\mathbf{c}_{t})\}_{t\in\mathcal{I}^{*}}$ and the target camera poses $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$ for the $M$ new frames to be generated and samples the new views:
@@
Specifically, we base our generator on the recently released pre-trained SEVA [46] model. Given the plug-and-play nature of our memory module, VMem can work with other image-set generators as well. While our model can operate with SEVA out-of-the-box, the original SEVA is trained with a large fixed total number of reference and target frames ($K+M=21$), which demands substantial computational resources that may be prohibitive for many applications. To address this, we fine-tune a more efficient version that operates with a reduced number of reference frames ($K=4$) and target views ($M=4$) and use it for our main experiments. Our ablation studies (Sec. 4.6) demonstrate that this fine-tuned version, when combined with VMem, achieves comparable performance to the original SEVA while delivering approximately 10$\times$ speedup at inference time.
Short-term ($50^{th}\text{frame}$) Long-term ($>200^{th}\text{frame}$)
LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
@@
VMem 0.287 18.49 0.406 17.12 0.219 0.039 0.493 13.12 0.183 27.15 0.811 0.499
Table 1: Quantitative single-view NVS comparison on RealEstate10K Lower is better for LPIPS, FID, $R_{\text{dist}}$, $T_{\text{dist}}$; higher is better for PSNR, SSIM. VMem surpasses all baseline methods on this benchmark, but its true advantage—spatial consistency—is not fully reflected, as the trajectories in RealEstate10K rarely revisit previously observed areas.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
@@
VMem 0.381 14.82 0.275 24.97 0.793 0.124
Table 2: Quantitative comparison on cycle trajectories from RealEstate10K. LPIPS (lower is better), PSNR (higher is better), $R_{\text{dist}}$ (lower is better), $T_{\text{dist}}$ (lower is better).
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
@@
VMem 0.472 14.11 0.121 1.204 0.387
Table 3: Quantitative comparison on cycle trajectories from Tanks and Temples. LPIPS (lower is better), PSNR (higher is better), SSIM (higher is better), $R_{\text{dist}}$ (lower is better), $T_{\text{dist}}$ (lower is better).
Context views $K$ View Retrieval LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $R_{\text{dist}}$downarrow $T_{\text{dist}}$downarrow
@@
Table 4: Ablation study on cycle trajectories from RealEstate10K. We compare different view retrieval strategies for two context view counts $K$. $K=17$ is the maximum context size in SEVA [46] for generating $M=4$ target views, while we use $K=4$ in our main experiments due to the $\sim$12$\times$ faster inference. The original SEVA strategy relies on camera distance for retrieval.
4 Experiments
We conduct experiments on long-term, short-term, and cycle-trajectory perpetual view generation tasks, comparing our models to previous open-source single-view approaches. Our results show that VMem can generate more coherent long-term scenes and demonstrates significantly better camera control than all existing open-source video generation methods.
Figure 6: Qualitative comparison of the cycle trajectory frames ($>400$ frames). Under the cycle trajectory setting, VMem can realistically infer and generate unseen content with significantly more precise camera control than both inpainting-based and multi-view/video-based baseline methods. Moreover, it retains memory of all previously generated frames, ensuring consistency when revisiting locations—an ability none of the multi-view/video-based baseline methods [42, 36, 23] possess.
4.1 Implementation Details
The novel view generator of VMem is trained on the RealEstate10K dataset [47] using SEVA [46] as a pretrained backbone. The number of context views $K$ is set to $4$, and the number of target views $M$ is also $4$. We randomly sample context views online during training. We use LoRA [14] with rank $256$ to fine-tune the base model. The model is trained for 600,000 iterations on 8 A40 GPUs, with a batch size of 24 per GPU. The model is optimized using AdamW with a learning rate of $3\times 10^{-6}$, a weight decay of $10^{-4}$, and a cosine annealing schedule. During inference, the classifier-free guidance scale is set to $3$. The scaling factor $\sigma$ for the point map is set to $0.03$, and $\alpha$ is set to $0.2$ in the surfel radius calculation.
4.2 Evaluation Metrics
We evaluate the quality of generated novel views based on three key factors: (1) the overall quality of generated novel views, (2) the model’s ability to preserve input image details across views, and (3) the alignment between generated camera poses and ground truth.
For (1), we compare the distribution of the generated novel views to that of the test set using Fréchet Image Distance (FID) [11]. For (2), we measure peak signal-to-noise ratio (PSNR) for pixel differences, along with LPIPS [45] and SSIM [35], as done in [23]. For (3), following [10], we express each estimated camera pose relative to the first frame and normalize the translation by the furthest frame. We use DUSt3R [34] to extract poses from generated views. We then calculate the rotation distance $R_{\text{dist}}$ by comparing the the ground truth and extracted rotation matrices of each generated novel view sequence, expressed as:
$R_{\text{dist}}=\arccos\left(\frac{\text{tr}(\mathbf{R}_{\text{gen}}\cdot\mathbf{R}^{T}_{\text{gt}})-1}{2}\right),$
where $\mathbf{R}$ denotes a rotation matrix. We also compute the translation distance $t_{\text{dist}}$ as:
4.3 Datasets and Evaluation Settings
We consider two real-world datasets for new view synthesis evaluation: RealEstate10K [47] and Tanks-and-Temples [15]. RealEstate10K comprises video clips of indoor scenes. We use the training split of RealEstate10K to train VMem. For evaluation, following [23, 41], we only consider test sequences with at least 200 frames. Starting from the first frame of the ground truth test sequence, we generate 20 images using 20 camera poses along the ground truth trajectory, spaced ten frames apart, with the final pose 200 frames from the initial view. Short-term evaluations use the fifth generated image, while long-term evaluations focus on the final image. For cycle evaluation, each model generates frames from the initial to the final camera pose and then back to the start, with evaluations conducted only on the return trajectory. The goal of this evaluation is to test the model’s ability to maintain scene consistency when revisiting previously traversed paths. For Tanks-and-Temples, which includes out-of-domain indoor and outdoor scenes featuring much larger motion speeds, we use all six advanced scenes and evaluate cycle-trajectory settings for the first 50 frames without skipping frames.
4.4 Cycle-Trajectory Perpetual View Generation
The key advantage of VMem is its ability to remember previously generated content across different locations, as demonstrated by the cycle-trajectory setting. VMem surpasses all baseline methods on both in-domain (Tab. 2) and out-of-domain (Tab. 3) benchmarks. Additionally, qualitative comparisons in Fig. 6 clearly illustrate that VMem can generate realistic content in unseen regions with significantly more precise camera control than both inpainting-based and multi-view/video-based baselines. Furthermore, VMem preserves the memory of previously generated frames, ensuring scene consistency upon revisiting previous locations—an ability that none of the multi-view/video-based baseline methods possess.
4.5 Long- and Short-Term Perpetual View Generation
To further validate our approach, we evaluate VMem on the widely recognized benchmark established in [23, 41] for both short-term and long-term perpetual generation tasks. As demonstrated in Tab. 1, VMem substantially outperforms all baseline methods across all metrics, achieving superior scene extrapolation and significantly improved camera control. These quantitative improvements are further supported by qualitative comparisons in Fig. 6. Notably, VMem enables the generation of consistent, high-quality long videos of scenes across diverse scenarios beyond those in RealEstate10K, as shown in Fig. 5. We believe this represents a crucial step forward in advancing the creation of immersive virtual worlds.
@@ 4.6 Ablation Study
We conduct an ablation study to assess the effectiveness of our proposed surfel-based memory indexing for context view retrieval in VMem. Specifically, we evaluate our approach in the cycle-trajectory setting and compare it against alternative view retrieval strategies: temporal-based, camera distance-based, and field of view-based retrieval. For temporal-based retrieval, we select the most recent $K$ views from memory. For camera distance-based retrieval, we select the top $K$ views with cameras closest to the target view camera—this strategy is also employed by SEVA [46], which serves as our backbone novel view generator. For field of view-based retrieval, we select the top $K$ views with the highest field of view overlap with the target view.
We also evaluate the performance trade-off between our lightweight, fine-tuned novel view generator using $K=4$ context views versus the original SEVA configuration with $K=17$ context views. The results are reported in Tab. 4.
Our findings demonstrate that VMem consistently improves the performance of the novel view generator across both context view configurations. The improvement is most pronounced when using fewer context views ($K=4$), highlighting the effectiveness of our surfel-based memory indexing.
Notably, our VMem with only $K=4$ views recovers most of the performance achieved with $K=17$ views in the original SEVA setting while delivering approximately 12×12$\times$12 × speed improvement. On an RTX 4090 GPU, our lightweight novel view generator takes 4.2 seconds per frame, while the original SEVA takes 50 seconds per frame.
@@ 5 Conclusion
We introduced Surfel-Indexed View Memory (VMem), a novel plug-and-play memory mechanism for long-term autoregressive scene generation using a surfel-indexed view memory. By anchoring past views to a surfel-based representation of the scene and retrieving the most relevant ones, our approach improves long-term scene consistency while reducing computational costs. Experiments on long-term scene synthesis benchmarks demonstrate that VMem outperforms existing methods in coherence and image quality. Our work advances scalable video generation with applications in virtual reality, gaming, and other domains.
@@ Acknowledgments.
Appendix A Average Pose Calculation
@@
Appendix B Autoregressive Point Map Prediction
@@
Appendix C Limitation and Discussion
@@
Since there is no established benchmark for evaluating long-term consistency in scene video generation, we adopt cyclic trajectories as a proxy for assessment. However, these trajectories remain relatively simple and contain only limited occlusions, which means the full potential of VMem in handling occlusions is not fully demonstrated. Moreover, existing evaluation metrics primarily capture low-level texture similarity in hallucinated content, rather than assessing true multi-view consistency—an inherent limitation of single-view perpetual generation. As such, there is a clear need for more standardized evaluation protocols, which we leave for future exploration.
@@
Due to limited computational resources, VMem was fine-tuned only on the RealEstate10K dataset [47], which primarily consists of indoor scenes and a limited number of outdoor real-estate scenarios. Consequently, the model may struggle to generalize well to broader contexts. Specifically, performance might degrade when dealing with natural landscapes or images containing moving objects, compared to indoor environments. However, we believe this limitation mainly stems from insufficient dataset diversity and computational constraints.
@@
Future Improvement.
Since our memory module relies heavily on the capabilities of the off-the-shelf image-set model and the point map predictor, the performance of VMem is expected to improve as these underlying models continue to advance.

### v3

@@ 3.2 Novel view generator
Figure 5: Long sequences with revisitations. We compare our VMem against a baseline without memory that relies solely on the last $K$ frames for context. Each sequence: input images (left), then generated images at selected frames. Our method (top rows) maintains consistency when revisiting observed regions, while the baseline (bottom rows) shows severe inconsistencies across extended sequences.
We use a camera-conditioned image-set generator $\psi$ to generate novel views. The generator $\psi$ takes $K$ retrieved reference views $\mathcal{V}^{*}=\{v_{t}=(\mathbf{x}_{t},\mathbf{c}_{t})\}_{t\in\mathcal{I}^{*}}$ and the target camera poses $\{\mathbf{c}_{T+m}\}_{m=1}^{M}$ for the $M$ new frames to be generated and samples the novel views:
@@
Specifically, we base our generator on the recent SEVA [46] model. Given the plug-and-play nature of our memory module, VMem can work with other image-set generators as well. We also fine-tune a more efficient version of SEVA that operates with a reduced number of reference frames. We provide more details in Sec. 4.1.
Short-term ($50^{th}\text{frame}$) Long-term ($\geq 200^{th}\text{frame}$)
LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$ $\downarrow$ $T_{\text{dist}}$ $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$ $\downarrow$ $T_{\text{dist}}$ $\downarrow$
@@
SEVA [46] ($K=17$) 0.293 18.33 0.382 17.29 0.223 0.118 0.455 13.98 0.216 23.82 1.125 0.742
VMem ($K=4$) 0.287 18.49 0.406 17.12 0.219 0.039 0.493 13.12 0.183 27.15 0.811 0.499
VMem ($K=17$) 0.293 18.33 0.382 17.29 0.223 0.118 0.452 14.09 0.227 23.56 0.982 0.432
Table 1: Single-view NVS on RealEstate10K. Frames are subsampled with 10-frame intervals. $K$ denotes the number of context views. In the short-term setting, all past frames fit in the context windows for both SEVA [46] and VMem; hence, SEVA [46] and VMem ($K=17$) yield identical results since VMem uses SEVA as the backbone. VMem ($K=4$) is fine-tuned on RealEstate10K for this setting, yielding better performance than SEVA [46] and VMem ($K=17$) in short-term scenarios. VMem surpasses baselines on most metrics in this benchmark. However, its true advantage—spatial consistency—is not fully reflected because RealEstate10K trajectories rarely revisit previously observed areas. We address this limitation in Sec. 4.3.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ $R_{\text{dist}}$ $\downarrow$ $T_{\text{dist}}$ $\downarrow$
@@
SEVA [46] ($K=17$) 0.401 11.82 0.217 24.72 0.902 0.492
VMem ($K=4$) 0.397 15.72 0.297 24.97 0.821 0.392
VMem ($K=17$) 0.304 18.15 0.377 24.18 0.892 0.165
Table 2: Results on cycle trajectories from RealEstate10K.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $R_{\text{dist}}$ $\downarrow$ $T_{\text{dist}}$ $\downarrow$
@@
VMem ($K=4$) 0.472 14.11 0.121 1.204 0.387
Table 3: Results on cycle trajectories from Tanks and Temples.
4 Experiments
We evaluate our method on established benchmarks for camera-conditioned autoregressive view generation from a single image, comparing our models to previous open-source approaches (Sec. 4.2). These benchmarks, however, have a critical limitation: camera trajectories rarely revisit previously observed regions, unlike real-world behavior where humans or robots often return to or look at previously explored areas. To address this limitation, we propose a cycle-trajectory evaluation protocol that extends these benchmarks by making the camera return to the starting position along the same path in reverse order (Sec. 4.3). Additionally, we collect a small dataset with trajectories that revisit observed regions for qualitative evaluation in Fig. 5. Finally, we conduct an ablation study in Sec. 4.4.
4.1 Experimental setup
Implementation details.
We use pre-trained SEVA [46] as the generator $\psi$. SEVA uses a fixed total number of reference and target frames ($K+M=21$). For our experiments, we use $M=4$ target views, which leaves $K=17$ context views. However, this configuration demands substantial computational resources that may be prohibitive for many applications. To address this limitation, we fine-tune a more efficient version with LoRA [13] on the RealEstate10K training split [47] that works with a reduced number of reference frames ($K=4$) and target views ($M=4$), and use this efficient version for our main experiments. We demonstrate (Secs. 4.3 and 4.4) that this fine-tuned version, when combined with VMem, achieves comparable performance to the original SEVA while delivering approximately $12\times$ speedup at inference time. We provide more technical details in Appendix A.
Datasets.
We consider two real-world datasets for novel view generation evaluation: RealEstate10K [47], comprising indoor scene video clips, and Tanks-and-Temples [14], including indoor and outdoor scenes with larger camera motions. For qualitative evaluation, we use in-the-wild images collected from the internet or captured with phone cameras as input views (see Fig. 5).
Evaluation metrics.
We evaluate generated novel view quality based on three factors: (1) overall quality of generated novel views, comparing generated novel view distributions to test set distributions using Fréchet Image Distance (FID) [10]; (2) the ability of the model to preserve image details across views, measuring peak signal-to-noise ratio (PSNR) for pixel differences, along with LPIPS [45] and SSIM [34], as in [22]; and (3) alignment between generated camera poses and ground truth, following [9], where we express estimated camera poses relative to the first frame and normalize translation by the furthest frame. We use DUSt3R [33] to extract poses from generated views. We calculate rotation distance $R_{\text{dist}}$ by comparing ground truth and extracted rotation matrices of each generated sequence:
$R_{\text{dist}}=\arccos\left(0.5\left(\text{tr}(\mathbf{R}_{\text{gen}}\mathbf{R}^{T}_{\text{gt}})-1\right)\right),$
where $\mathbf{R}$ denotes a rotation matrix. We compute translation distance $t_{\text{dist}}$ as:
Figure 6: Qualitative comparison of the cycle trajectory frames ($\geq 400$ frames). VMem autoregressively generates new views while maintaining memory of all previously generated frames, ensuring consistency when revisiting previously seen locations.
4.2 Short- and long-term view generation
We first evaluate VMem on the benchmark from [22, 42], which defines two scenarios: short-term and long-term novel view generation. Beginning from the first frame of each ground-truth test sequence, the model autoregressively generates images along the ground truth camera trajectories with 10-frame intervals. The short-term scenario evaluates the fifth generated image (50 frames from the initial view), while the long-term scenario evaluates the final image ($\geq$200 frames from the initial view). Following the protocol, we only consider samples with more than 200 frames for the long-term scenario and more than 50 frames for the short-term scenario. We report results in Tab. 1. VMem substantially outperforms all baselines on the key metrics, demonstrating superior scene extrapolation. However, the benchmark trajectories rarely revisit observed regions, limiting evaluation of long-term scene consistency.
Context
views $K$
View Retrieval LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ $R_{\text{dist}}$ $\downarrow$ $T_{\text{dist}}$ $\downarrow$
@@ 4.2 Short- and long-term view generation
Table 4: Ablation study on cycle trajectories from RealEstate10K. We compare different view retrieval strategies for two context view counts $K$
4.3 Long-term view generation with revisitations
To evaluate long-term scene consistency when revisiting observed regions—a scenario that rarely occurs in existing benchmarks—we propose a cycle-trajectory evaluation protocol as a proxy for real revisitation scenarios in human and robotic exploration. In this protocol, the model generates frames following trajectories from initial to final poses, then returns along the same path to the starting position. We evaluate metrics every 10 frames on the return trajectories, using all test sequences. Following this protocol, we report quantitative results in Tab. 2 and qualitative comparisons in Fig. 6. VMem outperforms all baselines across all metrics, demonstrating superior visual consistency when generating previously observed regions.
To validate generalization, we evaluate on Tanks-and-Temples containing indoor and outdoor scenes with more dynamic camera trajectories. We use all six advanced scenes and apply the same cycle-trajectory protocol using the first 50 frames, evaluating every frame without temporal subsampling. Results are in Tab. 3. VMem consistently outperforms all baselines across in-domain (Tab. 2) and out-of-domain (Tab. 3) evaluations on the majority of metrics.
We also qualitatively evaluate VMem across diverse in-the-wild scenes in Fig. 5, demonstrating generation of diverse, consistent, high-quality long videos.
@@ 4.4 Ablation study
We ablate our surfel-based memory indexing for context view retrieval in VMem. We evaluate our approach in the cycle-trajectory setting against three alternative retrieval strategies: (1) temporal-based retrieval, selecting the most recent $K$ views from memory, (2) camera distance-based retrieval, selecting top $K$ views with cameras closest to the target view camera—also used by SEVA [46], our backbone generator, and (3) field-of-view-based retrieval, selecting top $K$ views with highest field-of-view overlap with the target view. We also evaluate performance trade-offs between our lightweight, fine-tuned generator using $K=4$ context views versus original SEVA with $K=17$ context views. We report results in Tab. 4.
Our findings show VMem consistently improves generator performance across both numbers of context views. Improvement is most pronounced with fewer context views ($K=4$), highlighting the effectiveness of our surfel-based memory index.
Computational efficiency.
A key practical advantage of our approach is its computational efficiency. VMem with only $K=4$ views achieves $12\times$ speed improvement over the original SEVA while recovering most of the performance achieved with $K=17$ views. On an RTX 4090 GPU, our method generates frames in 4.2 seconds compared to 50 seconds for SEVA, representing a step towards real-time interactive scene exploration.
@@ 5 Conclusion
We introduced Surfel-Indexed View Memory (VMem), a novel plug-and-play memory module for long-term autoregressive scene generation using a surfel-indexed view memory. By anchoring past views to a surfel-based representation of the scene and retrieving the most relevant ones, our approach improves long-term scene consistency while reducing computational costs. Experiments on long-term scene synthesis benchmarks demonstrate that VMem outperforms existing methods in scene coherence while enabling the use of fewer context views, leading to significantly faster generation. Our work advances scalable video generation, with applications in virtual reality, gaming, and other domains.
@@ Acknowledgments.
Appendix A Implementation details
To address the computational cost of SEVA [46], which uses a large fixed total number of reference and target frames ($K+M=21$), we fine-tune a more efficient version that employs a reduced number of reference frames ($K=4$) and target views ($M=4$). We apply LoRA [13] with rank $256$ and randomly sample context views online during training. Training proceeds for 600,000 iterations on 8 A40 GPUs with a batch size of 24 per GPU, using the AdamW optimizer with a learning rate of $3\times 10^{-6}$, weight decay of $10^{-4}$, and a cosine annealing schedule. For inference, we set the classifier-free guidance scale to $3$, the point map scaling factor $\sigma$ to $0.03$, and $\alpha$ to $0.2$ for surfel radius calculation.
Appendix B Average pose calculation
@@
Appendix C Autoregressive point map prediction
@@
Appendix D Limitations and discussion
@@
Since there is no established benchmark for evaluating long-term consistency in scene video generation, we adopt cyclic trajectories as a proxy for assessment. However, these trajectories remain relatively simple and contain only limited occlusions, which means the full potential of VMem in handling occlusions is not fully demonstrated. Moreover, existing evaluation metrics primarily capture low-level texture similarity in hallucinated content, rather than assessing true multi-view consistency—an inherent limitation of single-view autoregressive generation. As such, there is a clear need for more standardized evaluation protocols, which we leave for future exploration.
@@
Due to limited computational resources, our more efficient version of the generator based on SEVA [46] was fine-tuned only on the RealEstate10K dataset [47]. This dataset primarily consists of indoor scenes and a limited number of outdoor real-estate scenarios. Consequently, the model may struggle to generalize to broader contexts, with performance potentially degrading when dealing with natural landscapes or images containing moving objects compared to indoor environments. We believe this limitation stems primarily from insufficient dataset diversity rather than fundamental model constraints.
@@
Future improvements.
Since our memory module relies heavily on the capabilities of the off-the-shelf image-set generator and the point map predictor, the performance of VMem is expected to improve as these underlying models continue to advance.